# Time Non-locality in Dark Matter and LSS

### v1

@@ 5.2 Connection with observation
We have found through Eq.(5.1) and (5.2), that the time non-local basis at third order contains one more operator than the time local basis. This implies that there exists one operator in the non-local basis that cannot be written in terms of the local basis. We have also shown that the time non-local basis is mathematically equivalent to the basis given in Eq.(5.3). The extra non-local operator, is easily extracted through the basis which is . The fact that time non-locality manifests itself at third order implies that it can be measured in one-loop power spectrum of through a contribution like or in the tree level four point function .
@@ 6 Discussion and Conclusion
Galaxy formation time is of order the evolution timescales of the Universe as a whole and hence is inherently a time non-local process. Indeed, in the context of the EFTofLSS, [1] has recently shown that time non-locality arises explicitly at the fifth order in bias expansion. However, a natural question is that if galaxy formation is intrinsically time-non-local in nature, why does this aspect show up only at the fifth order in the expansion? Our analysis reveals that delving beyond scalar observables that were considered in [1] can unveil non-locality even at lower orders. For instance, in the context of the effective stress tensor within the framework of EFTofLSS — — which transforms as a second-rank tensor under transformations, temporal non-locality manifests at the third order111111This non-locality in dark matter is anticipated due to the lack of separation between the time scales of long and short modes.. Another instance arises when incorporating selection operators in the bias expansion. These operators, acting as projections of tensorial objects along the line-of-sight, inherently possess the structures necessary to exhibit temporal non-locality at lower orders. Notably, in the case of Lyman- flux fluctuations, where the symmetry breaks down to SO(2) along the line of sight, temporal non-locality emerges again at the third order in the bias expansion121212Let us note that in [1] and in the current paper, we have only used SPT kernels which corresponds to non-renormalised matter overdensity and velocity divergence as given in (2.1).. This suggests that temporal non-locality might be observable in the 1-loop power spectrum for phenomena necessitating the inclusion of selection operators. This is in contrast to the bias expansion, where non-locality is only detectable at the two-loop power spectrum level for observables.
@@
In addition to exploring these physics concepts, our paper delved into several intriguing technical aspects. Typically, non-locality emerges at significantly higher orders in perturbative expansions. For instance, in the case of the bias expansion, it arises at the 5th order. Analyzing non-locality at these higher orders presents considerable technical challenges. Therefore, gaining insight into more efficient methods for investigating non-locality is crucial. In this study, we discovered that employing the basis [2] offers a more efficient approach to handling non-locality. Specifically, concerning biased tracers, we demonstrated that the correspondence between the basis and the time non-local basis extends up to the fifth order for invariant operators. It’s noteworthy that the system proves more effective in establishing the non-local basis, requiring only 52 scalar operators, a significant reduction compared to the approximately 130 scalar operators in the conventional basis. We have also established efficacy of basis when dealing with selection operators.
@@
In deriving these results, we have employed the usual method of writing selection operators in the bias expansion as given in Eq.(2.15). In this method we write all possible invariant operators at a given order. This is termed as the bottom-up approach to writing operators in the bias expansion. However, for the case of Lyman- forest, we can generate the same set of operators in the bias expansion of flux fluctuation () through renormalisation of the flux field. Due to the non-linear exponential map between the optical depth () and the total flux (), given as , contains operators with products of evaluated at the same point in space. This requires renormalisation of these composite operators. Taking the optical depth fluctuation in redshift space and renormalising the field, in principle, should generate all possible operators that appear in the bias expansion in Eq.(2.15). For more details see Appendix. B.
@@
In [62], it has been pointed out that renormalisation generates all operators that are allowed by the symmetries. It would be intriguing to check, whether renormalisation generates the time local or time non-local operators for the case of both and invariant operators.
@@
In order to derive the non-local basis for biased tracers, no assumption is made regarding whether tracers should follow continuity equation or not [1]. system, on the other hand, is derived by assuming that the tracers satisfy a continuity equation. However, we have shown that the non-local basis is related to the basis by an invertible map up to fifth for invariant operators and up to fourth order for invariant operators. It would be nice to have a physical understanding of this mathematical equivalence.
@@
Presence of time non-locality in the effective stress tensor as given in Table. 1, implies that the structure of kernels for dark matter overdensity as given in Eq.(2.1) should also change after renormalisation as these new non-local terms will be added as counterterms to the dark matter overdensity field [60]. As a result, if we write the operators in bias expansion with the renormalised kernels of EFT rather than using SPT kernels as given in Eq.(2.1), it would be interesting to see how the analysis of Secs. 4 and 5 will be modified. We leave this investigation for future exploration. Also in this study, we have not considered higher derivative operators in the bias expansion. It would be interesting to see how the results are modified when higher derivative operators are also considered for both and selection operators. Given the fact that time non-locality appears at fourth order in , which is a two derivative operator, one would expect the same for higher derivative operators. Exploring this aspect of bias expansion is left for future work.
@@
Recently, bootstrap ideas have been applied to write the kernels for the biased tracers [63]. Just using symmetry properties of the equations of motion, it has been shown that the kernels for biased tracers can be bootstrapped up to third order. It would be interesting to investigate time non-locality from a bootstrap perspective. Preliminary investigation along these lines suggests one interesting technical point, which is the presence of an even more efficient basis than the basis that we have employed in this paper.
@@
One can also do a full -body simulation and measure the contribution of time non-local operators found in this study and the studies preceding this [1]. An exhaustive -body simulation with generic initial conditions maybe computationally challenging. However, with simpler initial conditions, see for e.g. [64], such computations maybe tractable. Measuring the contributions due to time non-local operators in simulations will establish time non-locality as a generic feature of LSS.
@@ Appendix A Basis of vector operators () up to third order
Appendix B Comments on Renormalisation
@@
The total flux received depends on the optical depth () of neutral hydrogen along the line of sight of observation in an inverse manner. Optical depth is proportional to the number density of neutral hydrogen. So if the number density is high then there will be more absorption and hence flux will be reduced. The number density of neutral hydrogen is a direct tracer of the underlying dark matter. Being a direct tracer of the underlying dark matter, the optical depth which is proportional to the number density has the same bias expansion as given in Eq.(2.7). Mathematically, the total flux is related to the optical depth by an exponential map as follows [65, 66],
where is the mean optical depth and is the fluctuation in the optical depth in redshift space. Now, using Eq.(B.2), we can write how the flux fluctuation in the Lyman alpha field is related to the fluctuation in the optical depth which is given as [57],
@@
So, let’s look at the structure of operators that arise when we go to redshift space. Redshift space is related to the real space by a coordinate transformation given as [67, 68],
where and are the redshift space and real coordinates respectively and is the unit vector along the line of sight. Under the coordinate change given in Eq.(B.4) and imposing mass conservation, the optical depth fluctuation is expressed in momentum space as,
where has the same bias expansion as given in Eq.(2.7). We can expand Eq.(B.5) up to desired order in overdensity to get the bias expansion in redshift space. For example up to third order, we have,
where and we have used for brevity. From Eq.(B), we can see that a subset of selection operators are generated by going to redshift space.
We can do a time non-local expansion of Eq.(B) to get the set of non-local operators in redshift space. We have checked that up to third order, the non-local operators can be written in terms of the local operators. Hence, time non-locality does not appear in redshift space up to third order.
We can substitute the expression for from Eq.(B) in Eq.(B.3) to get the expression for flux fluctuation in redshift space. The second and third terms in Eq.(B.3) are composite operators. Therefore, it has been pointed out that the second and the third term in Eq.(B.3) need to be renormalised [57] appropriately. Upon renormalisation we expect that all operators allowed by the symmetry are generated as counterterms. Therefore, we get the same bias expansion as in Eq.(5.1). To generate all selection operators upto third order, we need third order tree level counterterms. The systematic method of renormalisation for local Eulerian halo biasing in the context of invariant operators is developed in [62]. From that study, it can be observed that to generate all counterterm at third order, we need to evaluate the quantity,
where are linear overdensity in dark matter. Now, has the expansion given in Eq.(B.3) which contains . The kernel for upto fifth order in momentum space is too big to deal with. Furthermore, identifying individual selection operators is impractical, therefore we leave that exercise for future and assume that all selection operators are generated when is renormalised to all loops.

### v2

@@ 5.2 Connection with observation
We have found through Eq.(5.1) and (5.2), that the time non-local basis at third order contains one more operator than the time local basis. This implies that there exists one operator in the non-local basis that cannot be written in terms of the local basis. We have also shown that the time non-local basis is mathematically equivalent to the basis given in Eq.(5.3). Among the operators given in Eq.(5.3), all operators except and are writable in terms of the time local basis given in Eq.(5.1). Interestingly, the linear combination is writable in terms of the local basis given in Eq.(5.1), revealing that this combination is time local. However, any other linear combination such as,
is not writable in terms of the time local basis and hence, is intrinsically time non-local. Following [Desjacques:2018pfv], we show that the time non-local operator given in Eq.(5.7) gives non trivial contribution to the 1-loop power spectrum.
Let us give a brief overview of the basic idea given in [Desjacques:2018pfv]. The power spectrum upto one loop has the following structure,
where is the linear growth factor given in Eq.(2.1). We will only be analysing the contribution coming from , as time non-locality manifests at third order. The general structure of this contribution has the following form,
where is one of the operators from the list given in Eq.(5.3). Any cubic operator can be expressed as,
Using Eq.(5.10) in Eq.(5.9) we get the most general structure of the 1-3 contribution,
Let us now look at the form of the possible kernels appearing in Eq.(5.11). One class of operators, that contains products of three , have kernel of the following form [Desjacques:2018pfv],
where is a projection operator made out of products of Kronecker delta and even powers of . Operators whose kernels take the form of Eq.(5.12) have been shown to be fully absorbed by counterterms, and therefore do not contribute independently to the 1-loop power spectrum [Desjacques:2018pfv]. Another class of kernels that appear at third order is given as,
where the projection operator is again made out of products of Kronecker delta and even powers of and in general, is different from the one given in Eq.(5.12). However, if it contains then Eq.(5.13) takes the same form as Eq.(5.12) and hence do not contribute to the 1-loop power spectrum. For more details see Appendix. C and D of [Desjacques:2018pfv].
The explicit form for both the operators given in Eq.(5.7) is as follows,
The last line of Eq.(5.2) is of the form given in Eq.(5.12), which implies that it can be absorbed by counterterms and does not contribute to the 1-loop power spectrum. However, all other terms in Eq.(5.2) give non-trivial contributions to the 1-loop power spectrum individually as well as in the linear combination given in Eq.(5.7). Hence, we conclude that the time non-local operator given in Eq.(5.7) is relevant at one-loop power spectrum. This suggests that time non-local effects can be measured in the power spectrum of at 1-loop through a contribution that goes like . At tree level, time non-locality will affect the four-point function as well.
@@ 6 Discussion and Conclusion
Galaxy formation time is of order the evolution timescales of the Universe as a whole and hence is inherently a time non-local process. Indeed, in the context of the EFTofLSS, [Donath:2023sav] has recently shown that time non-locality arises explicitly at the fifth order in bias expansion. However, a natural question is that if galaxy formation is intrinsically time-non-local in nature, why does this aspect show up only at the fifth order in the expansion? Our analysis reveals that delving beyond scalar observables, that were considered in [Donath:2023sav], can unveil non-locality even at lower orders. For instance, in the context of the effective stress tensor within the framework of EFTofLSS — — which transforms as a second-rank tensor under transformations, temporal non-locality manifests at the third order151515This non-locality in dark matter is anticipated due to the lack of separation between the time scales of long and short modes.. This implies that one necessarily has to use time non-local counterterms in the effective stress tensor beyond second order. However, in any calculation involving stress tensor only up to second order, one can simply use the local counterterms which are simpler to work with. Another instance where time non-locality is indispensable arises when incorporating selection operators in the bias expansion. These operators, acting as projections of tensorial objects along the line-of-sight, inherently possess the structures necessary to exhibit temporal non-locality at lower orders. Notably, in the case of Lyman- flux fluctuations, where the symmetry breaks down to SO(2) along the line of sight, temporal non-locality emerges again at the third order in the bias expansion161616Let us note that in [Donath:2023sav] and in the current paper, we have only used SPT kernels which corresponds to non-renormalised matter overdensity and velocity divergence as given in (2.1).. This suggests that temporal non-locality might be observable in the 1-loop power spectrum for phenomena necessitating the inclusion of selection operators. This is in contrast to the bias expansion, where non-locality is only detectable at the two-loop power spectrum level for observables. Therefore, we see that time non-locality manifests itself at different orders for different observables. Our analysis has pointed out this distinction.
@@
In addition to exploring these physics concepts, our paper delved into several intriguing technical aspects. Typically, non-locality emerges at significantly higher orders in perturbative expansions. For instance, in the case of the bias expansion, it arises at the fifth order. Analyzing non-locality at these higher orders presents considerable technical challenges. Therefore, gaining insight into more efficient methods for investigating non-locality is crucial. In this study, we discovered that employing the basis [Mirbabayi:2014zca] offers a more efficient approach to handling non-locality. The novel aspect of our study is to identify momentum structures in (time non-local) basis such as those given in Eq.(5.7), which contains inherently non-local-in-time physics. The method to generate different bases (time local, non-local and ), however, is not new and has been the object of many studies [Senatore:2014eva, Mirbabayi:2014zca, DAmico:2022ukl, Donath:2023sav, Abolhasani:2015mra, Vlah:2019byq]. Specifically, concerning biased tracers, we demonstrated that the correspondence between the basis and the time non-local basis extends up to the fifth order for invariant operators. It’s noteworthy that the system proves more effective in establishing the non-local basis, requiring only 52 scalar operators at fifth order to begin with, a significant reduction compared to the approximately 130 scalar operators in the conventional approach. We have also established the efficacy of the basis when dealing with selection operators.
@@
In deriving these results, we have employed the usual method of writing selection operators in the bias expansion as given in Eq.(2.14). In this method we write all possible invariant operators at a given order. This is termed as the bottom-up approach to writing operators in the bias expansion. However, for the case of Lyman- forest, we can generate the same set of operators in the bias expansion of flux fluctuation () through renormalisation of the flux field. Due to the non-linear exponential map between the optical depth () and the total flux (), given as , contains operators with products of evaluated at the same point in space. This requires renormalisation of these composite operators. Taking the optical depth fluctuation in redshift space and renormalising the field, in principle, should generate all possible operators that appear in the bias expansion in Eq.(2.14). For more details see Appendix. C.
@@
In [Assassi:2014fva], it has been pointed out that renormalisation generates all operators that are allowed by the symmetries. It would be intriguing to check, whether renormalisation generates the time local or time non-local operators for the case of both and invariant operators.
@@
In order to derive the non-local basis for biased tracers, no assumption is made regarding whether tracers should follow continuity equation or not [Donath:2023sav]. system, on the other hand, is derived by assuming that the tracers satisfy a continuity equation. However, we have shown that the non-local basis is related to the basis by an invertible map up to fifth for invariant operators and up to fourth order for invariant operators. It would be interesting to have a physical understanding of this mathematical equivalence.
One may wonder whether the equivalence of the and the time non-local basis extends beyond fifth order. Doing explicit computation of the basis and comparing will be very challenging as we know from fifth order. It would be nice to have a formal proof of this equivalence between the and the time non-local basis at all orders. There is already a hint to why this should happen given the fact that operators given in Eq.(2.16) are just Lagrangian space operators derived by Taylor expanding the Eulerian operators about the initial time slice [Mirbabayi:2014zca]. This Taylor expansion is similar to the Taylor expansion we do to obtain the time non-local operators as given in Eq.(3.7).
@@
Presence of time non-locality in the effective stress tensor as given in Table. 1, implies that the structure of kernels for dark matter overdensity as given in Eq.(2.1) should also change after renormalisation as these new non-local terms will be added as counterterms to the dark matter overdensity field [DAmico:2022ukl]. As a result, if we write the operators in bias expansion with the renormalised kernels of EFT rather than using SPT kernels as given in Eq.(2.1), it would be interesting to see how the analysis of Secs. 4 and 5 will be modified. We leave this investigation for future exploration. Also in this study, we have not considered higher derivative operators in the bias expansion. It would be interesting to see how the results are modified when higher derivative operators are also considered for both and selection operators. Given the fact that time non-locality appears at fourth order in , which is a two derivative operator, one would expect the same for higher derivative operators. Exploring this aspect of bias expansion is left for future work.
@@
Recently, bootstrap ideas have been applied to write the kernels for the biased tracers [DAmico:2021rdb]. Just using symmetry properties of the equations of motion, it has been shown that the kernels for biased tracers can be bootstrapped up to third order. It would be interesting to investigate time non-locality from a bootstrap perspective. Preliminary investigation along these lines suggests one interesting technical point, which is the presence of an even more efficient basis than the basis that we have employed in this paper.
@@
One can also do a full -body simulation and measure the contribution of time non-local operators found in this study and the studies preceding this [Donath:2023sav]. An exhaustive -body simulation with generic initial conditions maybe computationally challenging. However, with simpler initial conditions, see for e.g. [Karandikar:2023ozp], such computations maybe tractable. Measuring the contributions due to time non-local operators in simulations will establish time non-locality as a generic feature of LSS.
@@ Appendix A Basis of vector operators () up to third order
The extra term in the non-local basis coming in the will contribute in the the equation of motion given in Eq.(3.1) [DAmico:2022ukl]. Hence, the kernels for will change as compared to the case when the local vector operator is considered.
Appendix B Basis of scalar operators () up to fourth order
In Sec. 1, we have stated that for scalar operators (), time non-locality manifests at fourth order. To demonstrate the same, we explicitly state the set of local, non-local and basis for scalar operators upto fourth order. We will show that upto third order all systems form equivalent bases which are related by linear invertible maps. However, at fourth order, time non-local basis differs from time local basis and hence, no map exists between time local and non-local basis. Nevertheless, basis is related to time non-local basis at each order. Note that the operators appearing in are also relevant for higher derivative operators in galaxy bias.
Below, we list the independent set of operators at each order sequentially. At linear order in matter overdensity, there is only one independent operator in the local, non-local, as well as system,
| Local basis: | ||||
| Non-local basis: | ||||
where we have used the same convention for writing time non-local operators as used in Sec. 4.
Second Order
At second order, we get 4 independent scalar operators in the time local, non-local, and basis. The set of independent operators are as follows,
| Local basis: | ||||
| Non-local basis: | ||||
We have explicitly checked that all three bases given in Eq.(B) are related by linear maps to each other.
Third Order
At third order, as in the case of vector operators, we have 12 independent scalar operators in the local basis. In case of vector operators, time non-local and basis contains 13 operators as given in Eq.(A). However, in contrast to the case of vector operators, scalar operators in time non-local and basis contain 12 independent operators which are given as,
| Local basis: | ||||
| Non-local basis: | ||||
We have checked that the different bases given in Eq.(B) are related by invertible linear maps. Hence, there are no non-trivial consequences of time non-locality at third order for scalar operators. However, as we will see below, this will change at fourth order.
Fourth Order
At fourth order, there are 30 independent scalar operators in the time local basis, whereas, the time non-local and basis contains 31 independent operators. This implies that the time non-locality has non-trivial effects at fourth order. Below, we list the set of independent operators at time local, non-local and basis,
| Local basis: | ||||
| Non-local basis: | ||||
We have checked explicitly that time non-local and basis as given in Eq.(B) are related by an invertible linear map. The extra time non-local operator appearing Eq.(B) will affect the renormalised velocity divergence () as given in Eq.(3.1).
Appendix C Comments on Renormalisation
@@
The total flux received depends on the optical depth () of neutral hydrogen along the line of sight of observation in an inverse manner. Optical depth is proportional to the number density of neutral hydrogen. So if the number density is high then there will be more absorption and hence flux will be reduced. The number density of neutral hydrogen is a direct tracer of the underlying dark matter. Being a direct tracer of the underlying dark matter, the optical depth which is proportional to the number density has the same bias expansion as given in Eq.(2.6). Mathematically, the total flux is related to the optical depth by an exponential map as follows [Gunn:1965hd, Croft:1997jf],
where is the mean optical depth and is the fluctuation in the optical depth in redshift space. Now, using Eq.(C.2), we can write how the flux fluctuation in the Lyman alpha field is related to the fluctuation in the optical depth which is given as [Ivanov:2023yla],
@@
So, let’s look at the structure of operators that arise when we go to redshift space. Redshift space is related to the real space by a coordinate transformation given as [Matsubara:2007wj, Senatore:2014vja],
where and are the redshift space and real coordinates respectively and is the unit vector along the line of sight. Under the coordinate change given in Eq.(C.4) and imposing mass conservation, the optical depth fluctuation is expressed in momentum space as,
where has the same bias expansion as given in Eq.(2.6). We can expand Eq.(C.5) up to desired order in overdensity to get the bias expansion in redshift space. For example up to third order, we have,
where and we have used for brevity. From Eq.(C), we can see that a subset of selection operators are generated by going to redshift space.
We can do a time non-local expansion of Eq.(C) to get the set of non-local operators in redshift space. We have checked that up to third order, the non-local operators can be written in terms of the local operators. Hence, time non-locality does not appear in redshift space up to third order.
We can substitute the expression for from Eq.(C) in Eq.(C.3) to get the expression for flux fluctuation in redshift space. The second and third terms in Eq.(C.3) are composite operators. Therefore, it has been pointed out that the second and the third term in Eq.(C.3) need to be renormalised [Ivanov:2023yla] appropriately. Upon renormalisation we expect that all operators allowed by the symmetry are generated as counterterms. Therefore, we get the same bias expansion as in Eq.(5.1). To generate all selection operators upto third order, we need third order tree level counterterms. The systematic method of renormalisation for local Eulerian halo biasing in the context of invariant operators is developed in [Assassi:2014fva]. From that study, it can be observed that to generate all counterterm at third order, we need to evaluate the quantity,
where are linear overdensity in dark matter. Now, has the expansion given in Eq.(C.3) which contains . The kernel for upto fifth order in momentum space is too big to deal with. Furthermore, identifying individual selection operators is impractical, therefore we leave that exercise for future and assume that all selection operators are generated when is renormalised to all loops.